Instructor-created question Question Help * A factory manager is planning for the manufacture of plywood to be sold overseas. The fixed cost of operation is estimated at $800,000 per month while the variable cost is $155 per thousand board feet of plywood The selling price will depend on how much wil be produced and sold and is determined by board foet. Determine the range of profitable production the relationship, price per thousand board feet p- $600-0.05D, where D is the amount produced and sold in thousands of For prultable po deon prouchon must be at least thosands of board foet per month but not moro than□(Round to to nearest unt
Answer #1

The profitable production, production must be at least where the loss and profit is zero ( or breakeven)

Cost =800,000 + 155D

TR = P*D = 600D- 0.05D2

Now profit is zero or breakeven where Cost=TR

800,000 + 155D = 600D - 0.05D2

0.05D2 - 445D + 800,000 =0

solving this we get

D = 2500 or D= 6400

So,

for profitable production must be atleast 2500 thousands of board feet per month but not more than 6400.